# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of economic state variables that influence the distribution of asset returns. For instance, $\mathbf{Y}_{t}$ could include interest rates, macroeconomic indices, volatility factors, etc. We assume each component $Y_{i,t}$ follows an Ito diffusion:
@@
with $\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt.$ In vector form,
@@
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is a $k\times k$ diagonal matrix of diffusion coefficients (i.e., $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t}))$ where $\sigma_{Y,i}$ are the diffusion functions from the component-wise definition), and $\mathbf{W}_{t}^{Y}$ is a $k$-dimensional Brownian motion whose components may be correlated.
We consider a financial market with $n$ risky assets and one risk-free asset. The risk-free asset accumulates wealth at an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the current state. For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
$dS_{i,t}=S_{i,t}\bigl{(}\mu_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{X}_{i,t}\bigr{)}.$
Here, $\mu_{i}(t,\mathbf{Y}_{t})$ is the instantaneous expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ is the instantaneous volatility of asset $i$. The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$). In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions for the risky assets. The risky asset dynamics can be written compactly as
@@
Throughout, we work on a complete probability space $\bigl{(}\Omega,\mathcal{F},\mathbb{P}\bigr{)}$ that carries the $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. Let $\bm{\pi}_{t}$ be the $n\times 1$ vector representing the fraction of wealth invested in each risky asset at time $t$. The investor also chooses a consumption rate $C_{t}$. The dynamics of the investor’s wealth $X_{t}$ are governed by the following stochastic differential equation (SDE):
$dX_{t}=\Bigl{[}\,X_{t}\Bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}\Bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$ (1)
Here, $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns over the risk-free rate, and $\mathbf{1}$ is an $n\times 1$ vector of ones.
@@
Fix a finite horizon $T>0$. The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth dynamics $X_{t}$ via Eq. (1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions.
@@
    Measurability and continuity. Each map is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
@@
    Global Lipschitz continuity. There exists $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
@@
    $+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
@@
    Linear growth. There exists $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
@@
    $\leq K_{\mathrm{LG}}^{2}\bigl{(}1+\lVert\mathbf{y}\rVert^{2}\bigr{)},\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\bigl{(}1+\lVert\mathbf{y}\rVert\bigr{)}.$
Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the chosen controls $\bm{\pi}_{t},C_{t}$) admits a unique strong solution for sufficiently well-behaved controls; see Oksendal, (2013, Chs. 5-6) for details on existence and uniqueness theorems for SDEs.
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximises the expected discounted utility
@@
subject to the wealth dynamics (1) and the admissibility constraints stated later in Assumption 2. Here $\delta>0$ is the subjective discount rate and $\kappa\geq 0$ weights the utility of terminal wealth (a bequest motive).
Throughout we work with the Hyperbolic Absolute Risk Aversion (HARA) family of utilities,
@@
whose domain is the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$ with $x_{\min}:=-a/b$. HARA nests the familiar special cases CRRA $(a=0,b=1)$ and CARA $(\gamma=1,b=1)$. On its domain $U$ is $C^{2}$, strictly increasing and concave, and it satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$ (see Pennacchi, , 2008, §3.2).
We restrict our attention to control processes that are economically sensible and mathematically tractable.
@@
    Adaptedness. $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
@@
    Domain-admissible wealth. With $x_{\min}:=-a/b$ (see the HARA utility defined above) we require
@@
For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$. The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the matrix $\bm{\Omega}$:
$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ is the cross-covariance matrix.
Assumption 3 (Noise covariance).
The block covariance matrix $\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$ is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that $z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2}$ for all $z\in\mathbb{R}^{n+k}$. (Consequently, $\Psi$ and $\Phi$ are individually positive definite and hence invertible.)
Assumption 2(iii) gives the strict inequalities $0<C_{t}<\xi X_{t}$, while Assumption 2(iv) guarantees the wealth process stays inside the utility’s domain, $X_{t}>x_{\min}\;(=-a/b)$, almost surely. Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada limits on $(x_{\min},\infty)$, and because $\delta>0$, we obtain
$\mathbb{E}\!\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\!\bigl{[}U(X_{T})\bigr{]}<\infty,$
so the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
@@
where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]=\mathbb{E}[\cdot|X_{t}=x,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
Assumption 4 (Uniform non-degeneracy and utility).
  1. (a)
    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists $\lambda>0$ such that
    $\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\quad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
  2. (b)
    Utility specification. We use the HARA utility $U$ in (3); on its domain $x>x_{\min}$ it is $C^{2}$, strictly increasing ($U^{\prime}>0$), strictly concave ($U^{\prime\prime}<0$) and satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$.
  3. (c)
    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$, and the terminal-wealth weight satisfies $\kappa\geq 0$.
Part (a) ensures the diffusion is non-degenerate, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation. Part (b) specialises preferences to the HARA family; it reduces to CRRA when $a=0$ (irrespective of $b$), and to the familiar log-utility when $a=0$ and $\gamma=1$. Together with the coefficient regularity of Assumption 1, these conditions yield existence, uniqueness, and sufficient smoothness of the value function (e.g. $V\in C^{1,2,2}$); see (Oksendal, , 2013, Chs. 6–7).
In our numerical approach, we parameterize the control policies using neural networks (NNs). The portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, determined by the NN parameters $\theta$ and $\phi$:
$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
These parameters $\theta$ and $\phi$ are then optimized to maximize the objective $J$.

### v4

@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
We assume that each component $Y_{i,t}$ of the state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ evolves according to an Itô diffusion:
@@
where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying
$\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$
with constant correlations $\phi_{ij}$. In vector form, the state process satisfies
@@
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k}$ is the drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k\times k}$ is a diagonal matrix of diffusion coefficients—i.e.,
$\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t})),$
and $\mathbf{W}^{Y}_{t}\in\mathbb{R}^{k}$ is a vector of correlated Brownian motions.
We consider a financial market consisting of $n$ risky assets and one risk-free asset. The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the state variable. The price $S_{i,t}$ of the $i$-th risky asset evolves according to
$dS_{i,t}=S_{i,t}\left(\mu_{i}(t,\mathbf{Y}_{t})\,dt+\sigma_{i}(t,\mathbf{Y}_{t})\,dW^{X}_{i,t}\right),\quad i=1,\dots,n,$
where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility of asset $i$. The Brownian motions $W^{X}_{i,t}$ driving the risky asset dynamics satisfy
$\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt,\quad\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$
with constant correlations $\psi_{ij}$ and $\rho_{ij}$.
In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n}$ denote the vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n\times n}$ the diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}\in\mathbb{R}^{n}$ the Brownian motion vector for the risky assets. Then the dynamics of the vector of risky asset prices $\mathbf{S}_{t}\in\mathbb{R}^{n}$ can be compactly written as
@@
To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector
$\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}.$
The instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with
$\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-covariation between the two sources of risk.
Throughout, we work on a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. The evolution of the investor’s wealth is described by the following stochastic differential equation (SDE):
$dX_{t}=\left[\,X_{t}\left(r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\right)-C_{t}\,\right]dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t},$ (1)
where $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones. The term $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns relative to the risk-free rate.
@@
Fix a finite horizon $T>0$. The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth process $X_{t}$ via Eq. (1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions:
@@
    Measurability and continuity. Each coefficient is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
@@
    Global Lipschitz continuity. There exists a constant $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
@@
    $\quad+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
@@
    Linear growth. There exists a constant $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
@@
    $\qquad\leq K_{\mathrm{LG}}^{2}\left(1+\lVert\mathbf{y}\rVert^{2}\right),\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\left(1+\lVert\mathbf{y}\rVert\right).$
Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the control processes $\bm{\pi}_{t}$ and $C_{t}$) admits a unique strong solution for sufficiently well-behaved controls. [^5] See Oksendal, (2013, Chs. 5–6) for the existence and uniqueness of solutions to SDEs.
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximizes the expected discounted utility
@@
subject to the wealth dynamics in Eq. (1) and the admissibility conditions stated later in Assumption 2. Here, $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ governs the strength of the bequest motive by weighting terminal utility.
Throughout, we assume preferences are represented by a utility function from the Hyperbolic Absolute Risk Aversion (HARA) family:
@@
defined on the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$, with $x_{\min}:=-a/b$. To ensure that the utility function is well-defined and the wealth process $X_{t}$ remains strictly positive, we assume $x_{\min}\geq 0$ throughout. This restriction precludes bankruptcy and reflects standard modeling practice in continuous-time portfolio choice under HARA preferences. This class includes the Constant Relative Risk Aversion (CRRA) utility for $a=0,b=1$, and the Constant Absolute Risk Aversion (CARA) utility for $\gamma=1,b=1$.
On its domain, $U$ is twice continuously differentiable, strictly increasing, and strictly concave. It also satisfies the Inada conditions. [^6] See Pennacchi, (2008, §3.2) for further discussion of the properties required of utility functions.:
$U^{\prime}(x)\to\infty\text{ as }x\downarrow x_{\min},\qquad U^{\prime}(x)\to 0\text{ as }x\to\infty.$
We consider control processes that are both economically sensible and mathematically tractable.
@@
    Adaptedness. The process $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
@@
    Domain-admissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
@@
Assumption 3 (Risk covariance).
The block covariance matrix
$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$
is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that
$z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2},\qquad\forall z\in\mathbb{R}^{n+k}.$
As a consequence, both $\Psi$ and $\Phi$ are individually positive definite and thus invertible.
The assumptions are standard in continuous-time portfolio choice models. Specifically, Assumption 2(d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada conditions on $(x_{\min},\infty)$, and because $\delta>0$, we obtain the following integrability bounds:
$\mathbb{E}\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\left[U(X_{T})\right]<\infty.$
Therefore, the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable when starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
@@
where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]:=\mathbb{E}[\cdot\mid X_{t}=x,\,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation. The function $V$ is defined on $[0,T]\times(x_{\min},\infty)\times\mathbb{R}^{k}$.
Assumption 4 (Uniform non-degeneracy).
Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists a constant $\lambda>0$ such that
$\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\qquad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
Assumption 4 ensures the diffusion matrix in the wealth dynamics is uniformly positive definite, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation. We continue to work with the HARA utility function $U$ defined in Eq. (3), which is $C^{2}$, strictly increasing and concave on its domain $(x_{\min},\infty)$, and satisfies the Inada conditions. Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chs. 6–7).
In our numerical implementation, we parameterize the control policies using neural networks (NNs). Specifically, the portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, with functional forms determined by neural network parameters $\theta$ and $\phi$:
$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\qquad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
The parameters $\theta$ and $\phi$ are optimized to maximize the objective functional $J$ defined in Eq. (2).